1. The dean of a school of business is forecasting total student enrollment for this year (2019)'s summer session classes based on the following historical data:
YEAR TOTAL ENROLLMENT
  y
2015 2,000
2016 2,200
2017 2,800
2018 3,000
 
a) What is this year's forecast using a three-year simple moving average?
b) What is this year's forecast using a three-year weighted moving average with weights of 0.5, 0.3, and 0.22
c) What is this year's forecast using exponential smoothing with alpha=0.4, if year 2017's smoothed forecast was 2,600?
d) What is the slope (b) of the least squares trend line for these data?
e) What is the Y-intercept (a) of the least squares trend line for these data?
f) What is this year's forecast using the least squares trend line for these data?
Answer #1

(a) For Period t, let the actual enrollment be At and Forecast be Ft

For 3 year moving average, Ft = (At-1 + At-2 + At-3)/3

=> F2019 = (A2018 + A2017 + At2016)/3 = (3000 + 2800 + 2200)/3 = 2666.67

(b)

For 3 year weighted moving average, Ft = 0.5At-1 + 0.3At-2 + 0.2At-3

=> F2019 = 0.5A2018 + 0.3A2017 + 0.2At2016 = 0.5*3000 + 0.3*2800 + 0.2*2200 = 2780

(c)

Exponential forecasting coefficient ? = 0.40,

For Period t, Actual Demand is = At

Forecast = Ft

Exponential Forecast Ft = Ft-1 + ? (At-1 - Ft-1)

given, F2017 = 2600

A2017 = 2800

=> F2018 = F2017 + ? (A2017 - F2017) = 2600 + 0.4(2800 - 2600) = 2680

F2019 = F2018 + ? (A2018 - F2018) = 2680 + 0.4(3000 - 2680) = 2808

(d)

Let the year be denoted by x and enrollment by y

Let the Regression line be y = bo + b1x

where,

bo = ( Σy Σx2 - Σx Σxy ) / ( nΣx2 - (Σx)2 )

b1 = ( nΣxy - ΣxΣy ) / ( nΣx2 - (Σx)2 )

Year x Enrollment (y) x2 xy
2015 1 2000 1 2000
2016 2 2200 4 4400
2017 3 2800 9 8400
2018 4 3000 16 12000
Total 10 10000 30 26800

=> bo = ( 10000*30 - 10*26800 ) / ( 4*30 - 102 ) = 1600

b1 = ( 4*26800 - 10*10000 ) / ( 4*30 - 102 ) = 360

=> y = 1600 + 360x

Hence, slope = b1 = 360

(e) Y - intercept = bo = 1600

(f) For Year 2019, x = 5

=> y = 1600 + 360*5 = 3400